OpenStudy (anonymous):

A recipe for 1 lb of trail mix calls for 2.5 oz of chocolate pieces. How many ounces of chocolate pieces are needed to make 5 lb of trail mix? A. 0.5 oz B. 3.5 oz C. 7.5 oz D. 12.5 oz

OpenStudy (nonopro):

d

OpenStudy (anonymous):

i have more

OpenStudy (anonymous):

Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited 2/5 into a savings account. He deposited the rest into a checking account. How much money did Edwin deposit into his checking account?

OpenStudy (mathstudent55):

You can use a proportion: 1 lb is to 2.5 oz as 5 lb is to x \(\dfrac{1}{2.5} = \dfrac{5}{x} \) Cross multiply.

OpenStudy (anonymous):

12.5

OpenStudy (mathstudent55):

Correct. \(1x = 5 \times 2.5\) \(x = 12.5\)